OpenStudy (anonymous):

Factor k2 + 13k + 12.

OpenStudy (jiteshmeghwal9):

\[k^2+13k+12\] \[=k^2+12k+k+12\]\[=k(k+12)+1(k+12)\]\[=(k+1)(k+12)\]